OpenStudy (lovelyanna):

Which is a correct interpretation of the slope of the line in this graph? A. The volume of water in the tank is decreasing by 2.5 L per hour. B. The volume of water in the tank is decreasing by 2 L per hour. C. The volume of water in the tank is decreasing by 0.5 L per hour. D. The volume of water in the tank is decreasing by 0.4 L per hour.

OpenStudy (lovelyanna):

@jigglypuff314 ?

jigglypuff314 (jigglypuff314):

decreasing 1/2 liter per hour

OpenStudy (lovelyanna):

C, can you explain, please?

jigglypuff314 (jigglypuff314):

points (1, 9) and (5,7) 5-1=4 9-7=2 over a period of 4 hours, decreased by 2 2/4 = 1/2

OpenStudy (lovelyanna):

Oh, ok. So we just had to find the slope and simplify, and that's our answer!

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

If 5=x2, then is 7= y2?

jigglypuff314 (jigglypuff314):

huh?

OpenStudy (lovelyanna):

Hold on.. I think I'm getting messed up

OpenStudy (lovelyanna):

Can you tell me which is \[(x1, y1), (x2, y2)\]

jigglypuff314 (jigglypuff314):

oh yeah then you were right :) I got confused for a sec >,<

OpenStudy (lovelyanna):

I'm right about what?

OpenStudy (lovelyanna):

So the answer is not C?

jigglypuff314 (jigglypuff314):

answer is C and you were right about "If 5=x2, then is 7= y2"

OpenStudy (lovelyanna):

Oh, ok. So your sure its C? :P

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Ok, thanks :)

OpenStudy (lovelyanna):

I gtg now, thanks again for your help! Sorry I wasted your time

OpenStudy (lovelyanna):

Bye! :) ♥